OpenStudy (anonymous):

simplify:

OpenStudy (anonymous):

hold on, i'm trying to write the equation

OpenStudy (anonymous):

\[ \sqrt{\frac{5}{7}}\]

OpenStudy (anonymous):

\[7\sqrt{5}\]\[\frac{ \sqrt{5} }{ 7 }\]\[\frac{ \sqrt{35} }{ 5 }\]\[\frac{ \sqrt{35} }{ 7 }\]

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

7sqrt5?